OpenStudy (anonymous):

To what value does \[\sqrt{1+\sqrt{1+\sqrt{1+\sqrt{1+...}}}}\]converges?

OpenStudy (jamesj):

The back of the envelope way is to first assume it does converge to some value, call it x. In which case \[ x = \sqrt{1 + x} \] Then solve for x. But you also need to convince yourself that it does indeed converge so that this isn't' a nonsense calculation.

OpenStudy (anonymous):

seems to be 1.6180339887499

OpenStudy (jamesj):

Ugh. Closed form answer please!

OpenStudy (anonymous):

What's a closed form answer?

OpenStudy (anonymous):

oh you mean like an expression with a big sigma thing in it?

OpenStudy (jamesj):

you've written down an approximation of the answer. Completely unnecessary. You can write an exact, closed form of the solution.

OpenStudy (mr.math):

x=1/2(1+sqrt5) is the solution of x=sqrt(1+x).

OpenStudy (mr.math):

But how does this help?

OpenStudy (anonymous):

\[x = \frac{\sqrt{5}+1}{2}\]

OpenStudy (anonymous):

isn't that the golden ratio?

OpenStudy (anonymous):

why is it the golden ratio?

OpenStudy (jamesj):

There's a puzzle for you.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I don't even know what a golden ratio is.

OpenStudy (anonymous):

back of the envelop! i like it. actually hard part is showing it converges, easy part is solving a quadratic equation.

OpenStudy (anonymous):

you can google 'golden ratio' and get more information that you can possibly use.

OpenStudy (anonymous):

def fib(n): last = 0 this = 1 if n < 2: return n for i in xrange(1,n): new = last + this last = this this = new return this for i in xrange(1,101): print float(fib(i+1))/float(fib(i)) it also prints out the golden ratio

OpenStudy (anonymous):

i can't really read this but it looks like what you have is the limit of the quotients of the fibonnaci numbers, which is the golden ratio

OpenStudy (anonymous):

Hmm.... this all seems deja vu to me

OpenStudy (anonymous):

Excercise try to find X in here $$ X=5^0+ \sqrt{5^1+\sqrt{5^2+\sqrt{5^4+\sqrt{5^8+\sqrt{5^{16}+\sqrt{5^{32}+\dots}}}}}}$$

OpenStudy (jamesj):

That's a nice competition question.

OpenStudy (anonymous):

I think the solution can be implemented using python

OpenStudy (anonymous):

this should be in project euler

OpenStudy (jamesj):

It can be solved--and should be--with pencil and paper.

OpenStudy (anonymous):

alright. what's step 1?

OpenStudy (jamesj):

Think about it, given your original question. They're very related.

OpenStudy (anonymous):

alright what's the trick?

OpenStudy (anonymous):

It is easy when you know the trick.

OpenStudy (anonymous):

the trick is that the next number is equal to the previous number?

OpenStudy (anonymous):

or something?

OpenStudy (jamesj):

think about it for a while.

OpenStudy (anonymous):

is it the same answer but multiplied by 5?

OpenStudy (jamesj):

if it is, explain why. If not, what is it. Do some work.

OpenStudy (anonymous):

btw @JamesJ technically, you need to prove the convergence before writing $$ x = \sqrt{1 + x} $$ isn't ? ;)

OpenStudy (anonymous):

lol okay. I'll fire up google and if it doesn't help, I will think about it.

OpenStudy (jamesj):

@FFM. That is exactly what I said.

OpenStudy (anonymous):

By Google-ing you may be robing yourself of an opportunity to solve a good problem :)

OpenStudy (anonymous):

OH sorry I haven't read all of the discussions.

OpenStudy (anonymous):

is the answer 5?

OpenStudy (anonymous):

nice trick!

OpenStudy (anonymous):

$$ \frac{7+\sqrt{5}}{2}$$

OpenStudy (jamesj):

Yes, that's right.

OpenStudy (anonymous):

nooo you revealed the answer to me :(

OpenStudy (jamesj):

Now figure out why it's right.

OpenStudy (anonymous):

\[X = 1 + \sqrt{5 + X}\]

OpenStudy (anonymous):

almost there

OpenStudy (mr.math):

I computed the first few terms and found that (call the first summation Y): \[X=1+\sqrt{5}Y.\]

OpenStudy (anonymous):

it's been half an hour, and no luck for me :(

OpenStudy (mr.math):

\[\sqrt{5+\sqrt{5^2}}=\sqrt{5+5}=\sqrt{5}\sqrt{2}.\] \[\sqrt{5+\sqrt{5^2+\sqrt{5^4}}}=\sqrt{5+\sqrt{5^2+5^2}}=\sqrt{5+5\sqrt{2}}=\sqrt{5}\sqrt{1+\sqrt{2}}.\] \[\sqrt{5+\sqrt{5^2+\sqrt{5^4+\sqrt{5^8}}}}=\sqrt{5+\sqrt{5^2+\sqrt{5^4+5^4}}}\] \[=\sqrt{5+\sqrt{5^2+5^2\sqrt{2}}}=\sqrt{5+5\sqrt{1+\sqrt{2}}}=\sqrt{5}{\sqrt{1+\sqrt{1+\sqrt{2}}}}\]. Does this help?

OpenStudy (anonymous):

kind-of.... I still can't find that recurrence relation thing I'm supposed to look for

OpenStudy (mr.math):

I think you can see that\[X=5^0+\sqrt{5}Y.\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I still can't solve it

OpenStudy (mr.math):

\[X=1+\sqrt{5}.{\sqrt{5}+1 \over 2}=1+{5+\sqrt{5} \over 2}={7+\sqrt{5} \over 2}.\]

OpenStudy (mr.math):

I wouldn't be able to find it if James didn't do the first part though.

OpenStudy (anonymous):

*Agdgdgdgwngo feels enlightened.

OpenStudy (anonymous):

But how did you figure that trick out?

OpenStudy (anonymous):

*Agdgdgdgwngo is actually enlightened.

OpenStudy (anonymous):

so this is how we apply the golden ratio